Using the method of undetermined coefficients, find the general solution of the DE yiv -2ym+2yn=3e-x+2e-x+e-xsin x.


1
Expert's answer
2022-02-08T07:31:44-0500

"y^{(iv)} -2y'''+2y''=3e^{-x}+2e^{-x}+e^{-x}sin x=5e^{-x}+e^{-x}sin x"


characteristic equation:

"k^4-2k^3+2k^2=0"

"k^2(k^2-2k+2)=0"

"k_{1,2}=0"

"k_{3,4}=\frac{2\pm \sqrt{4-8}}{2}=1\pm i"


"y_c=c_1+c_2x+e^x(c_3cosx+c_4sinx)"


"y_{p1}=Ae^{-x}"

"Ae^{-x}+2Ae^{-x}+2Ae^{-x}=5e^{-x}"

"A=1"

"y_{p1}=e^{-x}"


"y_{p2}=e^{-x}(Acosx+Bsinx)"

"y'_{p2}=-e^{-x}(Acosx+Bsinx)+e^{-x}(-Asinx+Bcosx)="

"=e^{-x}((B-A)cosx-(A+B)sinx)"

"y''_{p2}=-e^{-x}((B-A)cosx-(A+B)sinx)+e^{-x}((A-B)sinx-(A+B)cosx)="

"=e^{-x}(-2Bcosx+2Asinx)"

"y'''_{p2}=-2e^{-x}(Asinx-Bcosx)+2e^{-x}(Acosx+Bsinx)="

"=2e^{-x}((A+B)cosx+(B-A)sinx)"

"y^{(iv)}_{p2}=-2e^{-x}((A+B)cosx+(B-A)sinx)+"

"+2e^{-x}(-(A+B)sinx+(B-A)cosx)=-4e^{-x}(Acosx+Bsinx)"


"-4e^{-x}(Acosx+Bsinx)-4e^{-x}((A+B)cosx+(B-A)sinx)+"

"+2e^{-x}(-2Bcosx+2Asinx)=e^{-x}sin x"


"-4A-4(A+B)-4B=0\implies A+B=0"

"-4B-4(B-A)+4A=1\implies A-B=1"

"A=1\/2,B=-1\/2"

"y_{p2}=e^{-x}(cosx-sinx)\/2"


"y=y_c+y_{p1}+y_{p2}="

"=c_1+c_2x+e^x(c_3cosx+c_4sinx)+e^{-x}+e^{-x}(cosx-sinx)\/2"